# The equivalence of asymptotic average shadowing and vague specification properties and its consequences

### v1

@@ 3 Pseudo-orbits for $(X,T)$
Let $(X,T)$ be a topological dynamical system. A sequence $\underline{x}\in X^{\infty}$ is an asymptotic average pseudo-orbit for $(X,T)$ if and only if $\underline{x}$ is a vague pseudo-orbit for $(X,T)$.
@@
Observe that for each $n\in\mathbb{N}_{0}$ so that $S^{n}(\underline{x})\in{\mathcal{U}}$ there exists $z_{n}\in X$ such that $\rho(x_{n},z_{n})<\delta$ and $\rho(x_{n+1},T(z_{n}))<\delta$. Thus, by the triangle inequality, we obtain
@@
For the converse, let $\underline{x}\in X^{\infty}$ be an asymptotic average pseudo-orbit for $(X,T)$. Fix $\varepsilon>0$ and $k\in\mathbb{N}$. Set
@@
4 Tracing properties
In the same way that the original definition of pseudo-orbit motivated generalizations, the classical shadowing property inspired alternative types of tracing properties.
The average shadowing property introduced by Blank [3] is one of the alternative notions.
@@ 4 Tracing properties
We say that a topological dynamical system $(X,T)$ has the average shadowing property if for every $\varepsilon>0$ there exists $\delta>0$ such that for every $\delta$-average pseudo-orbit $\underline{x}\in X^{\infty}$ for $T$ there exists $z\in X$ such that
@@
In [14], Kulczycki, Kwietinak and Oprocha provided an equivalent formulation for the average shadowing property for the case where the topological dynamical system is chain mixing.
Definition 4.2.
It is not hard to see that $\delta$-pseudo-orbits are also $\delta$-average pseudo-orbits. Roughly speaking, Kulczycki, Kwietniak and Oprocha showed that it is enough to assure that pseudo-orbits can be traced on average.
Theorem 4.3 (Theorem 3.6 in [14]).
@@
Inspired by the works of Blank [3], and Eirola, Nevanlinna and Pilyugin [8], Gu introduced the asymptotic average shadowing property in [9].
@@
We say that a topological dynamical system $(X,T)$ has the asymptotic average shadowing property if for every asymptotic average pseudo-orbit $\underline{x}\in X^{\infty}$ for $(X,T)$ there exists $z\in X$ such that
@@
The average shadowing property is known to be a consequence of the asymptotic average shadowing property (see [14, 22]). One can see that the converse is true provided that $(X,T)$ is surjective and Besicovitch complete.
@@
By Corollary 4.2 in [22], the asymptotic average shadowing property implies the average shadowing property. Moreover, by Corollary 30 in [15], if $(X,T)$ has the asymptotic average shadowing property, then $(X,T)$ is Besicovitch complete.
Conversely, assume that $(X,T)$ is Besicovitch complete and has the average shadowing property. Let $\underline{x}\in X^{\infty}$ be an asymptotic average pseudo-orbit. By Proposition 3.7, there exists an asymptotic pseudo-orbit $\underline{y}\in X^{\infty}$ such that
@@
It follows from Proposition 3.4 that $\underline{y}$ is a $\delta$-average pseudo-orbit for any $\delta>0$. Therefore, by the average shadowing property, for every $\varepsilon>0$ there exists a point $z^{(\varepsilon)}\in X$ such that $\rho_{B}(\underline{z}_{T}^{(\varepsilon)},\underline{y})<\varepsilon.$
@@
We now focus on another tracing property, namely, the vague specification property, introduced by Kamae in [11].
@@
We say that a topological dynamical system $(X,T)$ has the vague specification property if for every vague pseudo-orbit $\underline{x}\in X^{\infty}$ for $(X,T)$ there is $z\in X$ such that for every $\varepsilon>0$ the following holds:
@@
If we combine Propositions 2.3 and 2.2, we see that the notions of tracing used in Definitions 4.4 and 4.6 are actually equivalent. As a consequence of Theorem 3.12, we obtain the equivalence between Definitions 4.4 and 4.6.
@@
We first assume that $(X,T)$ has the vague specification property. Fix an asymptotic average pseudo-orbit $\underline{x}\in X^{\infty}$. By Theorem 3.12, $\underline{x}$ is also a vague pseudo-orbit. Hence, by the vague specification property, there exists $z\in X$ such that for every $\varepsilon>0$ it holds
@@
Conversely, assume that $(X,T)$ has the asymptotic average shadowing property and take $\underline{x}\in X^{\infty}$ which is a vague pseudo-orbit. By Theorem 3.12, $\underline{x}\in X^{\infty}$ is also an asymptotic average pseudo-orbit for $(X,T)$. Since $(X,T)$ has the asymptotic average shadowing property, there exists $z\in X$ such that $\rho_{B}(\underline{z}_{T},\underline{x})=0$. Applying Lemmas 2.2 and 2.3, we obtain
@@
5 Tracing properties on shift spaces
@@
For every $n\in\mathbb{N}_{0}$, we denote by $\mathcal{L}_{n}(X)\subset\mathscr{A}^{*}$ the set of all words $w\in\mathcal{L}(X)$ such that $|w|=n$. Hence, it holds
@@
We say that a shift space $X$ is mixing if for every pair of words $u,v\in\mathcal{L}(X)$ there exists $N\in\mathbb{N}$ such that for every $n\geq N$ there is a word $w\in\mathcal{L}_{n}(X)$ such that $uwv\in\mathcal{L}(X)$.
@@
We denote an oriented $\mathscr{A}$-labelled (multi) graph by a triple $G=\left(V,E,\tau\right)$, where $V$ is a set of vertices, $E$ is a set of edges, and a label map $\tau\colon E\to\mathscr{A}$. Every $e\in E$ has two end points, an initial vertex $i(e)\in V$ and terminal vertex $t(e)\in V$. Let $\ell\in\mathbb{N}$. A sequence of edges $\{e_{1},\ldots,e_{\ell}\}\subset E$ is called a path of length $\ell$ if there exists a sequence of vertices $\{v_{1},\ldots,v_{\ell}\}\subset V$ with $i(e_{i})=v_{i}$, and $t(e_{i})=v_{i+1}$ for every $1\leq i<\ell$.
We define the shift $X_{G}$ to be the set of all sequences $x\in\mathscr{A}^{\infty}$ such that $x$ is the sequence of labels of an infinite path on a graph $G$ and their shifts. A shift space $X$ is called sofic if there exists a labelled graph $G=\left(V,E,\tau\right)$ such that the set of vertices $V$ is finite and $X=X_{G}$.
@@
We say that $w\in\mathcal{L}(X)$ is a synchronizing word for a shift space $X$ if $uw,wv\in\mathcal{L}(X)$ implies $uwv\in\mathcal{L}(X)$. A shift space $X$ is called synchronized if it is transitive and has a synchronizing word. In the particular case of hereditary shifts, it is not hard to see that we can assume the synchronizing word to be composed only by 00’s. More precisely, we have the following lemma whose proof is left to the reader.
@@
The original specification property introduced by Bowen [4] for general topological dynamical systems has an equivalent formulation for shift spaces.
@@
We remark that there are mixing shift spaces without the specification property (see [10, 16]).
In 1988, Bertrand [1] proved that shift spaces with specification always have synchronizing words.
Theorem 5.6 (Theorem 1 in [1]).
@@
We will discuss convergence of sequences of points in $\mathscr{A}^{\infty}$ with respect to the $\bar{d}$ pseudometric. We remark that since $\bar{d}$ is a pseudometric, a limit point, if exists, might not be unique. We define the notion of Cauchy sequence and completeness with respect to the $\bar{d}$ pseudometric in the natural way.
@@
We say that a sequence $\{z^{(n)}\}_{n=1}^{\infty}\subset X$ is $\bar{d}$-Cauchy if for every $\varepsilon>0$ there exists $N\in\mathbb{N}$ such that for every $n,m\geq N$ it holds
@@
By the uniform equivalence between the Besicovitch and $\bar{d}$ pseudometrics (Proposition 5.7), it follows from Corollary 30 in [15] that the asymptotic average shadowing property is a sufficient condition for $\bar{d}$-completeness.
@@
In the particular case of shift spaces, a notion of tracing called $\bar{d}$-shadowing property, which is closely related to the average shadowing property, was introduced by Konieczny, Kupsa and Kwietniak in [13].
@@
As mentioned in [13], a shift space with the average shadowing property also possesses the $\bar{d}$-shadowing property. Moreover, if we assume surjectivity of the shift space, then the $\bar{d}$-shadowing property implies the average shadowing property. Due to the lack of proofs for these statements, for the sake of completeness, we will prove them.
@@
Set $E=\bigcup_{i\geq 2}(n_{i}-m,n_{i}]$. Note that if $n\in\mathbb{N}_{0}\backslash E$, then for its correspondent $i$ as above, it holds
@@
It remains to prove that $\underline{x}\in X^{\infty}$ is a $\delta$-average pseudo-orbit. Note that, also by construction, for every $n\in\mathbb{N}_{0}\backslash E$ it holds $x^{(n)}_{[1,m]}=x^{(n+1)}_{[0,m)}$, that is,
@@
Let $0<\delta\leq\frac{\varepsilon^{\prime}}{2}$ be such that every $\delta$-average pseudo-orbit is $\frac{\varepsilon^{\prime}}{2}$-traced on average by a point in $X$ and $N\in\mathbb{N}$ be provided by Lemma 5.13. Fix a sequence of words $\{w^{(i)}\}_{i=1}^{\infty}\subset\mathcal{L}(X)$ satisfying $|w^{(i)}|\geq N$ for every $i\in\mathbb{N}$.
By Lemma 5.13, there exists a $\delta$-average pseudo-orbit $\underline{x}\in X^{\infty}$ so that $\rho_{B}(\underline{w}_{\sigma},\underline{x})<\delta$. The average shadowing property yields the existence $z\in X$ such that $\rho_{B}(\underline{x},\underline{z}_{\sigma})<\frac{\varepsilon^{\prime}}{2}$. Therefore, by the triangle inequality, we obtain $\rho_{B}(\underline{w}_{\sigma},\underline{z}_{\sigma})<\varepsilon^{\prime}$. Finally, by the choice of $\varepsilon^{\prime}$, we conclude that $\bar{d}(w,z)<\varepsilon$. ∎
Our proof for a converse of Theorem 5.14 requires the shift space to be chain mixing. Lemma 9 in [13] guarantees that, in our context, surjectivity is sufficient.
Lemma 5.15 (Lemma 9 in [13]).
@@
Let $m\in\mathbb{N}$ be such that $\frac{N}{m}<\frac{\varepsilon}{2}$ and $0<\delta<\frac{\varepsilon}{2}$ be such that $\delta\leq 2^{-m}$. Fix a $\delta$-pseudo-orbit $\underline{x}\in X^{\infty}$. Note that $\delta\leq 2^{-m}$ implies $x^{(n)}_{[1,m]}=x^{(n+1)}_{[0,m)}$ for every $n\in\mathbb{N}_{0}$. As a consequence, for every $k\in[0,m)$ it holds
@@
We will use an argument similar to the one in Lemma 5.13 to prove that $\rho_{B}(\underline{w}_{\sigma},\underline{x})<\delta$. Set $E=\cup_{i\in\mathbb{N}}(im-N,im]$. It follows from (8) that for every $n\in\mathbb{N}_{0}\backslash E$ it holds $\rho(\sigma^{n}(w),x^{(n)})\leq 2^{-N}$. Note that $\bar{d}(E)\leq\frac{N}{m}$, thus $\rho_{B}(\underline{w}_{\sigma},\underline{x})<\frac{\varepsilon}{2}.$
@@
Since the $\bar{d}$-shadowing and the average shadowing properties are equivalent for surjective shift spaces, the equivalence of $\bar{d}$ and Besicovitch pseudometrics, together with Theorem 4.5, yields through the following corollary.

### v2

@@ 3 Pseudo-orbits in topological dynamical systems
Let $(X,T)$ be a topological dynamical system. A sequence $\underline{x}\in X^{\infty}$ is an asymptotic average pseudo-orbit in $(X,T)$ if and only if $\underline{x}$ is a vague pseudo-orbit in $(X,T)$.
@@
Observe that for each $n\in\mathbb{N}_{0}$ such that $S^{n}(\underline{x})\in{\mathcal{U}}$ there exists $z_{n}\in X$ such that $\rho(x_{n},z_{n})<\delta$ and $\rho(x_{n+1},T(z_{n}))<\delta$. Thus, by the triangle inequality, we obtain
@@
For the converse, let $\underline{x}\in X^{\infty}$ be an asymptotic average pseudo-orbit in $(X,T)$. Fix $\varepsilon>0$ and $k\in\mathbb{N}$. Set
@@
Figure 1: illustration of the triangle inequality.
4 Tracing properties of topological dynamical systems
In the same way as the original definition of pseudo-orbit motivated generalizations, the classical shadowing property inspired alternative types of tracing properties.
The average shadowing property introduced by Blank [Bla88] is one of the alternative notions.
@@ 4 Tracing properties of topological dynamical systems
We say that a topological dynamical system $(X,T)$ has the average shadowing property if for every $\varepsilon>0$ there exists $\delta>0$ such that for every $\delta$-average pseudo-orbit $\underline{x}\in X^{\infty}$ in $(X,T)$ there exists $z\in X$ such that
@@
In [KKO14], Kulczycki, Kwietniak and Oprocha provided an equivalent formulation for the average shadowing property for the case where the topological dynamical system is chain mixing. It is not hard to see that $\delta$-pseudo-orbits are also $\delta$-average pseudo-orbits, and roughly speaking, Kulczycki, Kwietniak and Oprocha showed that it is enough to assure that pseudo-orbits can be traced on average.
Theorem 4.2 (Theorem 3.6 in [KKO14]).
@@
Inspired by the works of Blank [Bla88], and Eirola, Nevanlinna and Pilyugin [ENP97], Gu introduced the asymptotic average shadowing property in [Gu07].
@@
We say that a topological dynamical system $(X,T)$ has the asymptotic average shadowing property if for every asymptotic average pseudo-orbit $\underline{x}\in X^{\infty}$ in $(X,T)$ there exists $z\in X$ such that
@@
The average shadowing property is known to be a consequence of the asymptotic average shadowing property (see [KKO14, WOC16]). One can see that the converse is true provided that $(X,T)$ is surjective and Besicovitch complete.
@@
By Corollary 4.2 in [WOC16], the asymptotic average shadowing property implies the average shadowing property. Moreover, by Corollary 30 in [KŁ17], if $(X,T)$ has the asymptotic average shadowing property, then $(X,T)$ is Besicovitch complete.
Conversely, assume that $(X,T)$ is Besicovitch complete and has the average shadowing property. Let $\underline{x}\in X^{\infty}$ be an asymptotic average pseudo-orbit. By Lemma 3.1 in [KKO14] $(X,T)$ is chain mixing, hence, by Proposition 3.8, there exists an asymptotic pseudo-orbit $\underline{y}\in X^{\infty}$ such that
@@
It follows from Proposition 3.4 that $\underline{y}$ is a $\delta$-average pseudo-orbit in any $\delta>0$. Therefore, by the average shadowing property, for every $\varepsilon>0$ there exists a point $z^{(\varepsilon)}\in X$ such that $\rho_{B}(\underline{z}_{T}^{(\varepsilon)},\underline{y})<\varepsilon.$
@@
Finally, we focus on the vague specification property, introduced by Kamae in [Kam76].
@@
We say that a topological dynamical system $(X,T)$ has the vague specification property if for every vague pseudo-orbit $\underline{x}\in X^{\infty}$ in $(X,T)$ there is $z\in X$ such that for every $\varepsilon>0$ the following holds:
@@
If we combine Propositions 2.3 and 2.2, we see that the notions of tracing used in Definitions 4.3 and 4.5 are actually equivalent. As a consequence of Theorem 3.13, we obtain the equivalence between Definitions 4.3 and 4.5.
@@
We first assume that $(X,T)$ has the vague specification property. Fix an asymptotic average pseudo-orbit $\underline{x}\in X^{\infty}$. By Theorem 3.13, $\underline{x}$ is also a vague pseudo-orbit. Hence, by the vague specification property, there exists $z\in X$ such that for every $\varepsilon>0$ it holds that
@@
Conversely, assume that $(X,T)$ has the asymptotic average shadowing property and take $\underline{x}\in X^{\infty}$ which is a vague pseudo-orbit. By Theorem 3.13, $\underline{x}\in X^{\infty}$ is also an asymptotic average pseudo-orbit in $(X,T)$. Since $(X,T)$ has the asymptotic average shadowing property, there exists $z\in X$ such that $\rho_{B}(\underline{z}_{T},\underline{x})=0$. Applying Lemmas 2.2 and 2.3, we obtain
@@
5 Tracing properties of shift spaces
@@
For every $n\in\mathbb{N}_{0}$, we denote by $\mathcal{L}_{n}(X)\subset\mathscr{A}^{*}$ the set of all words $w\in\mathcal{L}(X)$ such that $|w|=n$. Hence, the following holds
@@
We say that a shift space $X$ is transitive if for every pair of words $u,v\in\mathcal{L}(X)$ there exists $w\in\mathcal{L}(X)$ such that $uwv\in\mathcal{L}(X)$. We say that a shift space $X$ is mixing if for every pair of words $u,v\in\mathcal{L}(X)$ there exists $N\in\mathbb{N}$ such that for every $n\geq N$ there is a word $w\in\mathcal{L}_{n}(X)$ such that $uwv\in\mathcal{L}(X)$.
@@
We denote an oriented $\mathscr{A}$-labelled (multi) graph by a triple $G=\left(V,E,\tau\right)$, where $V$ is a set of vertices, $E$ is a set of edges, and $\tau\colon E\to\mathscr{A}$ is a so-called label map. Every $e\in E$ has two end points, an initial vertex $i(e)\in V$ and terminal vertex $t(e)\in V$. Let $\ell\in\mathbb{N}$. A sequence of edges $\{e_{1},\ldots,e_{\ell}\}\subset E$ is called a path of length $\ell$ if there exists a sequence of vertices $\{v_{1},\ldots,v_{\ell}\}\subset V$ with $i(e_{i})=v_{i}$, and $t(e_{i})=v_{i+1}$ for every $1\leq i<\ell$.
We define the shift $X_{G}$ to be the set of all sequences $x\in\mathscr{A}^{\infty}$ such that $x$ is the sequence of labels of an infinite path on a graph $G$. A shift space $X$ is called sofic if there exists a labelled graph $G=\left(V,E,\tau\right)$ such that the set of vertices $V$ is finite and $X=X_{G}$.
@@
We say that $w\in\mathcal{L}(X)$ is a synchronizing word for a shift space $X$ if $uw,wv\in\mathcal{L}(X)$ implies $uwv\in\mathcal{L}(X)$. A shift space $X$ is called synchronized if it is transitive and has a synchronizing word. In the particular case of hereditary shifts, it is not hard to see that we can assume the synchronizing word to be composed only of 00s. More precisely, we have the following lemma whose proof is left to the reader.
@@
The original specification property introduced by Bowen [Bow71] for general topological dynamical systems has an equivalent formulation for shift spaces.
@@
We remark that there are mixing shift spaces without the specification property (see [Jun11, KŁO16]).
In 1988, Bertrand [Ber88] proved that shift spaces with specification always have synchronizing words.
Theorem 5.6 (Theorem 1 in [Ber88]).
@@
We will discuss convergence of sequences of points in $\mathscr{A}^{\infty}$ with respect to the $\bar{d}$ pseudometric. We remark that since $\bar{d}$ is a pseudometric, a limit point, if it exists, might not be unique. We define the notion of Cauchy sequence and completeness with respect to the $\bar{d}$ pseudometric in the natural way.
@@
We say that a sequence $\{z^{(n)}\}_{n=1}^{\infty}\subset X$ is $\bar{d}$-Cauchy if for every $\varepsilon>0$ there exists $N\in\mathbb{N}$ such that for every $n,m\geq N$ we have
@@
By the uniform equivalence between the Besicovitch and $\bar{d}$ pseudometrics (Proposition 5.7), it follows from Corollary 30 in [KŁ17] that the asymptotic average shadowing property is a sufficient condition for $\bar{d}$-completeness.
@@
In the particular case of shift spaces, a notion of tracing called $\bar{d}$-shadowing property, which is closely related to the average shadowing property, was introduced by Konieczny, Kupsa and Kwietniak in [KKK23].
@@
As mentioned in [KKK23], a shift space with the average shadowing property also possesses the $\bar{d}$-shadowing property. Moreover, if we assume surjectivity of the shift space, then the $\bar{d}$-shadowing property implies the average shadowing property. Due to the lack of proofs for these statements, for the sake of completeness, we will prove them.
@@
Set $E=\bigcup_{i\geq 2}(n_{i}-m,n_{i}]$. Note that if $n\in\mathbb{N}_{0}\backslash E$, then for its correspondent $i$ as above, it holds that
@@
It remains to prove that $\underline{x}\in X^{\infty}$ is a $\delta$-average pseudo-orbit. Note that, also by construction, for every $n\in\mathbb{N}_{0}\backslash E$ we have $x^{(n)}_{[1,m]}=x^{(n+1)}_{[0,m)}$, that is,
@@
Let $0<\delta\leq\frac{\varepsilon^{\prime}}{2}$ be such that every $\delta$-average pseudo-orbit is $\frac{\varepsilon^{\prime}}{2}$-traced on average by a point in $X$, and let $N\in\mathbb{N}$ be provided by Lemma 5.13. Fix a sequence of words $\{w^{(i)}\}_{i=1}^{\infty}\subset\mathcal{L}(X)$ satisfying $|w^{(i)}|\geq N$ for every $i\in\mathbb{N}$.
By Lemma 5.13, there exists a $\delta$-average pseudo-orbit $\underline{x}\in X^{\infty}$ such that $\rho_{B}(\underline{w}_{\sigma},\underline{x})<\delta$. The average shadowing property yields the existence $z\in X$ such that $\rho_{B}(\underline{x},\underline{z}_{\sigma})<\frac{\varepsilon^{\prime}}{2}$. Therefore, by the triangle inequality, we obtain $\rho_{B}(\underline{w}_{\sigma},\underline{z}_{\sigma})<\varepsilon^{\prime}$. Finally, by the choice of $\varepsilon^{\prime}$, we conclude that $\bar{d}(w,z)<\varepsilon$. ∎
Our proof of a converse of Theorem 5.14 requires the shift space to be chain mixing. Lemma 9 in [KKK23] guarantees that, in our context, surjectivity is sufficient.
Lemma 5.15 (Lemma 9 in [KKK23]).
@@
Let $m\in\mathbb{N}$ be such that $\frac{N}{m}<\frac{\varepsilon}{2}$ and $0<\delta<\frac{\varepsilon}{2}$ be such that $\delta\leq 2^{-m}$. Fix a $\delta$-pseudo-orbit $\underline{x}\in X^{\infty}$. Note that $\delta\leq 2^{-m}$ implies $x^{(n)}_{[1,m]}=x^{(n+1)}_{[0,m)}$ for every $n\in\mathbb{N}_{0}$. As a consequence, for every $k\in[0,m)$ the following holds
@@
We will use an argument similar to the one in Lemma 5.13 to prove that $\rho_{B}(\underline{w}_{\sigma},\underline{x})<\delta$. Set $E=\cup_{i\in\mathbb{N}}(im-N,im]$. It follows from (9) that for every $n\in\mathbb{N}_{0}\backslash E$ it holds that $\rho(\sigma^{n}(w),x^{(n)})\leq 2^{-N}$. Note that $\bar{d}(E)\leq\frac{N}{m}$, thus $\rho_{B}(\underline{w}_{\sigma},\underline{x})<\frac{\varepsilon}{2}.$
@@
Since the $\bar{d}$-shadowing and the average shadowing properties are equivalent for surjective shift spaces, the equivalence of $\bar{d}$ and Besicovitch pseudometrics, together with Theorem 4.4, yields the following corollary.